# Phase structure of holographic superconductors in an Einstein-scalar-Gauss-Bonnet theory with spontaneous scalarization

### v1

@@ Abstract
Holographic superconductor phase transition and spontaneous scalarization are triggered by the instability of the underlying vacuum black hole spacetime. Although both hairy black hole solutions are closely associated with the tachyonic instability of the scalar degree of freedom, they are understood to be driven by distinct causes. It is, therefore, interesting to explore the interplay between the two phenomena in the context of a scenario where both mechanisms are present. To this end, we investigate the Einstein-scalar-Gauss-Bonnet theory in asymptotically anti-de Sitter spacetime with a Maxwell field. On the one hand, the presence of the charged scalar and Maxwell fields in anti-de Sitter spacetime furnishes the celebrated framework for a holographic superconductor. On the other hand, the non-minimal Gauss-Bonnet coupling between the scalar field and the gravitational sector triggers spontaneous scalarization. We show that these different origins for the tachyonic mass behave independently and can be recognized clearly by the distinctive natures of their effective potentials. However, near the transition curve, the holographic superconductor and spontaneous scalarization are found to be largely indistinguishable regarding both the radial profile and effective potential. Moreover, the employed numerical scheme, as well as the boundary conditions, are essentially identical. This raises the question of whether the hairy black holes triggered by different mechanisms are smoothly joined by a phase transition or whether these are actually identical solutions. To assess the transition more closely, we evaluate the phase diagram in terms of temperature and chemical potential. We discovered a smooth but first-order transition between the two hairy solutions by explicitly evaluating Gibbs free energy and its derivatives. In particular, one can elaborate a thermodynamic process through which a superconducting black hole transits into a scalarized one by raising or decreasing the temperature. Also, as the underlying gravitational system attains the solution that minimizes the Gibbs free energy, a scalarized black hole essentially overwhelms the superconducting one. Exhausting the underlying phase space, we analyze the properties and the interplay between the two hairy solutions. The holographic superconductor phase is observed to flip over to the other side of the transition curve when it drops below the critical temperature at vanishing Gauss-Bonnet coupling. Possible implications of the present study are addressed.
@@ I Introduction
One pertinent topic in black hole physics concerns a series of “no-hair” theorems and their evasion [10, 11, 12]. These theorems state that all the information about a black hole is determined by its mass, charge, and angular momentum. On the other hand, substantial insights can also obtained by exploring the scenarios when the prior condition of such theorems becomes invalid. The latter might give rise to hairy black hole solutions owing to various mechanisms [13, 14, 15]. The celebrated holographic superconductor is primarily owing to the tachyonic instability in the asymptotically anti-de Sitter (AdS) spacetime, complemented with the presence of a Maxwell field and a charged scalar [16, 17]. The asymptotic AdS spacetime is crucial to evading the prerequisite of no-hair theorems in asymptotically flat spacetimes [18]. More recently, spontaneous scalarization has aroused much attention in the literature. In its original form, it refers to the scenario where the scalar degree of freedom is nonminimally coupled to the Gauss-Bonnet curvature [19, 20, 21], giving rise to the emergence of hairy black hole solutions. Subsequently, it was observed that black hole spontaneous scalarization may take place in a more general context where the scalar field is nonminimallly coupled to source terms furnished by various types of matter fields, inclusively the Maxwell invariant [22], Chern-Simons invariant [23], and the Ricci scalar [24]. Further developments on spontaneous scalarization involve spin-induced [25, 26, 27, 28], nonlinear [29, 30, 31, 32], and dynamical descalarization scenarios [33, 34, 35, 36, 37, 38].
@@
As discussed in [39], it is essential to note that the tachyonic instability, being a sufficient condition, does not always align with the onset of marginally stable quasinormal modes. Nonetheless, the instability of the bald black hole solution is ascertained by the unstable quasinormal modes, while the subsequent transition to a hairy black hole can be confirmed by explicitly deriving the non-vanishing bound state solution [20, 41, 42] and evaluating the entropies of the scalarized black hole and compare it against that of its bald counterpart [19]. Given that the occurrence of unstable quasinormal modes is a weaker condition whose onset does not always warrant tachyonic instability, it was argued in [40] that spontaneous scalarization is caused by the Gregory-Laflamme instability [44]. In other words, another mechanism might trigger scalarization before the effective potential eventually becomes a potential well.
@@ II Einstein-scalar-Gauss-Bonnet Model
$<O_{i}>=\sqrt{2}\psi_{i},\ \ i=1,2$ (17)
with the boundary condition $\epsilon_{ij}\psi_{j}=0$. The asymptotical behavior of Eq. (14) can be used to extract the density and chemical potential.
@@ III Two types of hairy black hole solutions
In this section, we demonstrate that two distinct types of hairy black hole solutions associated with, respectively, holographic superconductor and scalarization, coexist in the model. When presented in terms of temperature and Gauss-Bonnet coupling, the phase diagram will be evaluated and is presented in Fig. 1. The system comprises three phases: thermalized vacuum represented by a phase of bald black hole solutions, holographic superconductor, and scalarization phases indicated by hairy black hole solutions. As further elaborated below, to identify and explore the transition between the two hairy black hole phases, one may present the phase diagram regarding temperature and chemical potential, shown in Fig. 2. Moreover, the thermodynamic properties of their interpretation in the dual field theory are then elaborated, and in particular, the free energy of obtained spacetime configurations are evaluated. We relegate the stability analysis of the underlying bald black hole solutions to Sec. III.3.
Figure 1: The phase diagram shown in terms of temperature $\tilde{T}$ and Gauss-Bonnet coupling $\lambda$. The blue region represents the holographic superconductor phase, and the solid black curve indicates the boundary where the transition to a bald black hole occurs. The red region corresponds to the spontaneous scalarization phase, so the black dashed line represents the critical value of the Gauss-Bonnet coupling. The calculations have been carried out by adopting $r_{h}=q=L=1$ and $m_{e}^{2}=-2$.
@@ III.1 Numerical procedure
For simplicity, the numerical calculations are carried out using $r_{h}=q=L=1$ and setting effective mass to $m^{2}_{e}=-2$, above the Breitenlohner-Freedman bound for stability. By comparing the obtained radial profiles $\psi(r)$ and $\phi(r)$ to the asymptotical form Eq.(15), the values of $\mu$ and $\rho$ are extracted. The resulting family of solutions is characterized by two variables, namely, the coupling $\lambda$ and $\psi(r_{h})$. One then employs the system’s scaling invariance by using Eqs. (18)-(21) to cover the remainder of the parameter space.
@@ III.2 Transitions among the black holes
Using the AdS/CFT dictionary, one can reiterate the phase diagram shown in Fig. 1 in terms of intensive quantities, namely, temperature and chemical potential. The results are presented in Fig. 2. When comparing Fig. 2 against Fig. 1, one observes a few intriguing features of the phase structure of the system.
The solid black curve bridges the transition between the holographically superconducting black hole and a hairless one, as shown in the left panel of Fig. 2. The latter corresponds to the region of elevated temperature and lower chemical potential. A dashed black curve indicates the transition between the holographic superconductor and spontaneous scalarization, which will be elaborated further by evaluating the free energy. However, the spontaneous scalarization phase does not extend and occupies the remainder of the phase space. It can be verified numerically that the temperature and chemical potential remain finite even at the limit $\lambda\to+\infty$. Also, a background hairless RN black hole spans the entire parameter space.
Figure 5: Chemical potential as a function of the Gauss-Bonnet coupling in the high temperature (left) and low temperature (right) regions. The blue curves represent the holographic superconductors while the red curves describe the spontaneous scalarization, and the vertical gray dashed line is the critical coupling $\lambda_{c}$.
@@
The results in the left and middle panels imply that the two phases successively flip to the other side of the transition curve in a narrow region of the phase diagram, as shown in the right panel of Fig. 2. Such an intriguing phenomenon can be attributed to the non-monotonic behavior of chemical potential as a function of Gauss-Bonnet coupling at different temperatures. At an elevated temperature related to the left panel of Fig. 2, the chemical potential decreases as the coupling constant increases. Conversely, at a lower temperature corresponding to the middle panel of Fig. 2, the chemical potential increases monotonically with increasing coupling. The above results are explicitly shown in the left and right panels of Fig. 5.
@@
Figure 6: The free energy and its first-order derivative concerning the temperature as functions of temperature for given chemical potential. The blue curves represent the holographic superconductors while the red curves describe the spontaneous scalarization, and the vertical gray dashed line indicates the transition point.
@@ III.3 Instability analysis using effective potential
In this subsection, we complement our analysis by studying the tachyonic instability of the effective potentials. On the one hand, for a holographic superconductor, it is understood that the coupling between the scalar and Maxwell field in asymptotic AdS spacetime leads to tachyonic instability [17]. On the other hand, the coupling between the scalar field and the Gauss-Bonnet curvature also gives rise to tachyonic instability [20, 41, 42], and it has been argued that the origin of the instability is Gregory-Laflamme type [40]. This section delves into the effective potential Eq. (11) of the scalar perturbations regarding the underlying instabilities. In literature, the tachyonic instability is primarily attributed to a negative effective mass extracted from examining the master equation of scalar perturbations. For a holographic superconductor, a negative effective scalar mass is attributed to the condensation of the Maxwell field in AdS spacetime. For spontaneous scalarization, this is due to the contribution coming from the non-vanishing coupling between the scalar and higher curvature term. The probe limit is considered for simplicity, and the asymptotic form of the Maxwell field is adopted for numerical calculations.
@@
As pointed out in [40], the effective potential of tachyonic instability in AdS spacetime is featured by a positive barrier near the event horizon that smoothly converges to a given value at infinity. In the left panel of Fig. 8, the resulting effective potentials of a pure holographic superconductor agree with such a behavior. In the present model, this can be achieved by assuming $\lambda=0$. The different curves in the plot are obtained by taking different values for the charge $q$. A positive potential barrier is formed near the event horizon for all the cases. An increase of the charge $q$ causes the potential barrier to dissipate more rapidly as the radial coordinate increases. The negative values of the effective potential are attributed to the condensation of the Maxwell field, giving rise to tachyonic instability in scalar perturbations. Also, a more significant coupling between the scalar and Maxwell fields leads to a higher transition temperature for holographic superconductivity [17].
@@
Figure 8: The calculated effective potentials for different model parameters. Left: The effective potential for the holographic superconductors with $\lambda=0$, evaluated for different charges $q$. Right: The effective potential for the spontaneously scalarized black hole with $q=0$, evaluated for different values of $\lambda$.
@@
Conversely, a pure spontaneously scalarized black hole metric can be obtained in the present model by taking $q=0$. The corresponding effective potentials are evaluated and shown in the right panel of Fig. 8, where one varies the Gauss-Bonnet coupling $\lambda$. One observes that the obtained effective potential’s main feature differs from that of a holographic superconductor. In particular, a potential well is formed near the event horizon, and the depth of this well gradually increases as the Gauss-Bonnet coupling $\lambda$ grows. Therefore, the potential well close to the horizon caused by the Gauss-Bonnet coupling is primarily understood to cause the tachyonic instability. Such a characteristic of the effective potential has been extensively discussed in the literature of spontaneous scalarization [20, 39], which was attributed to the Gregory-Laflamme instability in Ref. [40]. The above discussions further confirm the distinct nature of the holographic superconductivity and spontaneous scalarization phases, which are substantially triggered by entirely distinct mechanisms. Besides the thermodynamic quantities, such as the free energy, the difference is also manifested by the effective potential.

### v3

@@ Abstract
Holographic superconductor phase transition and spontaneous scalarization are triggered by the instability of the underlying vacuum black hole spacetime. Although both hairy black hole solutions are closely associated with the tachyonic instability of the scalar degree of freedom, they are understood to be driven by distinct causes. It is, therefore, interesting to explore the interplay between the two phenomena in the context of a scenario where both mechanisms are present. To this end, we investigate the Einstein-scalar-Gauss-Bonnet theory in asymptotically anti-de Sitter spacetime with the presence of a Maxwell field. Even though different origins for the tachyonic mass behave independently and can be recognized by the distinctive natures of their effective potentials, it is shown that near the transition curve, the holographic superconductor and spontaneous scalarization are found to be largely indistinguishable. This raises the question of whether the hairy black holes triggered by different mechanisms are smoothly joined by a phase transition or whether these are actually identical solutions. To assess the transition more closely, we evaluate the phase diagram in terms of temperature and chemical potential and discover a smooth but first-order transition between the two hairy solutions by explicitly evaluating Gibbs free energy and its derivatives. In particular, one can elaborate a thermodynamic process through which a superconducting black hole transits into a scalarized one by raising or decreasing the temperature. Exhausting the underlying phase space, we analyze the properties and the interplay between the two hairy solutions.
@@ I Introduction
One pertinent topic in black hole physics concerns a series of “no-hair” theorems and their evasion [10, 11, 12]. These theorems state that all the information about a black hole is determined by its mass, charge, and angular momentum. On the other hand, substantial insights can also obtained by exploring the scenarios when the prior condition of such theorems becomes invalid. The latter might give rise to hairy black hole solutions owing to various mechanisms [13, 14, 15]. The celebrated holographic superconductor is primarily owing to the tachyonic instability in the asymptotically anti-de Sitter (AdS) spacetime, complemented with the presence of a Maxwell field and a charged scalar [16, 17]. In this framework, the asymptotic AdS spacetime is crucial to evading the prerequisite of no-hair theorems in asymptotically flat spacetimes [18].
More recently, an alternative mechanism for hairy black holes, known as black hole spontaneous scalarization, has been proposed. In its original form, it refers to the scenario where the scalar degree of freedom is nonminimally coupled to the Gauss-Bonnet curvature in an asymptotically Minkowski spacetime, giving rise to the emergence of hairy black hole solutions [19, 20, 21]. In this regard, one pivotal feature of the theory is its ability to evade the no-hair theorem in asymptotically flat spacetimes. In the literature, the notion of black hole spontaneous scalarization quickly garnered significant attention. It was observed that black hole spontaneous scalarization may take place in a more general context where the scalar field is nonminimallly coupled to source terms furnished by various types of matter fields, inclusively the Maxwell invariant [22], Chern-Simons invariant [23], and the Ricci scalar [24]. Further developments on spontaneous scalarization involve spin-induced [25, 26, 27, 28], nonlinear [29, 30, 31, 32], and dynamical descalarization scenarios [33, 34, 35, 36, 37, 38]. This mechanism is also extendable to cases with a cosmological constant [39, 40, 41, 42, 43, 44, 45], contributing to its prominence. A comprehensive survey of recent progress can be found in the review [46].
@@
As discussed in [47], it is also essential to note that the tachyonic instability, being a sufficient condition, does not always align with the onset of marginally stable quasinormal modes. Nonetheless, the instability of the bald black hole solution is ascertained by the unstable quasinormal modes, while the subsequent transition to a hairy black hole can be confirmed by explicitly deriving the non-vanishing bound state solution [20, 49, 50] and evaluating the entropies of the scalarized black hole and compare it against that of its bald counterpart [19]. Given that the occurrence of unstable quasinormal modes is a weaker condition whose onset does not always warrant tachyonic instability, it was argued in [48] that spontaneous scalarization is caused by the Gregory-Laflamme instability [58]. In other words, another mechanism might trigger scalarization before the effective potential eventually becomes a potential well.
@@ II Einstein-scalar-Gauss-Bonnet Model
$<O_{1}>=\sqrt{2}\psi_{1}.$ (17)
The asymptotical behavior of Eq. (14) can be used to extract the density and chemical potential.
@@ III Two types of hairy black hole solutions
In this section, we demonstrate that two distinct types of hairy black hole solutions associated with, respectively, holographic superconductor and scalarization, coexist in the model. Using the numerical scheme presented in Sec. III.1, the phase diagram will be evaluated and presented in terms of temperature and Gauss-Bonnet coupling shown in Fig. 1. The system comprises three phases: thermalized vacuum represented by a phase of bald black hole solutions, holographic superconductor, and scalarization phases indicated by hairy black hole solutions. As will be validated in Sec. III.2, to identify and explore the transition between the two hairy black hole phases, one may also present the phase diagram in terms of temperature and chemical potential, shown in Fig. 2. We note that the latter two thermodynamic quantities are independent ones not constrained by scaling laws, while they facilitate the study of phase transition’s order. Moreover, the thermodynamic properties of their interpretation in the dual field theory are then elaborated, and in particular, the free energy of obtained spacetime configurations are evaluated. We relegate the stability analysis of the underlying bald black hole solutions to Sec. III.3.
Figure 1: The phase diagram shown in terms of temperature $\tilde{T}$ and Gauss-Bonnet coupling $\lambda$. The blue region represents the holographic superconductor phase, and the solid black curve indicates the boundary where the transition to a bald black hole occurs. The red region corresponds to the spontaneous scalarization phase, so the black dashed line represents the critical value of the Gauss-Bonnet coupling. The calculations have been carried out by adopting $r_{h}=L=1$, $q=1$ and $m_{e}^{2}=-2$.
@@ III.1 Numerical procedure
For simplicity, the numerical calculations are carried out using $r_{h}=L=1$ and $q=1$. In particular, following [59, 17, 62, 63, 64], we set the effective mass to $m^{2}_{e}=-2$, which is above the Breitenlohner-Freedman bound for stability. By comparing the obtained radial profiles $\psi(r)$ and $\phi(r)$ to the asymptotical form Eq.(15), the values of $\mu$ and $\rho$ are extracted. The resulting family of solutions is characterized by two variables, namely, the coupling $\lambda$ and $\psi(r_{h})$. One then employs the system’s scaling invariance by using Eqs. (18)-(21) to cover the remainder of the parameter space.
@@ III.2 Transitions among the black holes
The phase diagram presented in Fig. 1 is closely related to the numerical employed scheme, which is standard in the literature. However, to explore the underlying phase transition between the two hairy black holes, it is meaningful to show the phase diagram in terms of thermodynamic variables such as temperature and chemical potential. Using the AdS/CFT dictionary, one can reiterate the phase diagram shown in Fig. 1 in terms of intensive quantities, namely, temperature and chemical potential. The results are presented in Fig. 2. It is noted that the chemical potential does not remain constant when the system evolves along a vertical line with given $\lambda$, as shown in the bottom row of Fig. 5. When comparing Fig. 2 against Fig. 1, one observes a few intriguing features of the phase structure of the system.
The solid black curve bridges the transition between the holographically superconducting black hole and a hairless one, as shown in the left panel of Fig. 2. The latter corresponds to the region of elevated temperature and lower chemical potential. A dashed black curve indicates the transition between the holographic superconductor and spontaneous scalarization, which will be elaborated further by evaluating the free energy. In this region, one may consider the following process with given chemical potential for an initially thermalized bald black hole with elevated temperature. As the temperature gradually decreases, the scalar field will condensate and form a holographic supercondutor via a second-order phase transition by tranversing the solid black curve. Subsequently, as the temperature further decreases, the system transits into a scalarized black hole through a first-order transition by crossing the dashed black curve. However, the above qualitative properties regarding the phase division does not change as one goes to higher-temperature region. In other words, even though the solid black curve and dashed black one asymptotically approach each other in Fig. 1 at the high-temperature limit, the two-phase transitions do not actually intersect in the phase space presented in terms of temperature and chemical potential. Moreover, the spontaneous scalarization phase does not extend and occupies the remainder of the phase space. It can be verified numerically that the temperature and chemical potential remain finite even at the limit $\lambda\to+\infty$. Also, a background hairless RN black hole spans the entire parameter space.
Figure 5: The top row shows the chemical potential as a function of the Gauss-Bonnet coupling in the high temperature (top-left) and low temperature (top-right) regions, where the blue curves represent the holographic superconductors while the red curves describe the spontaneous scalarization, and the vertical gray dashed line is the critical coupling $\lambda_{c}$. The bottom row shows the chemical potential as a function of temperature for different Gauss-Bonnet couplings.
@@
The results in the left and middle panels imply that the two phases successively flip to the other side of the transition curve in a narrow region of the phase diagram, as shown in the right panel of Fig. 2. Such an intriguing phenomenon can be attributed to the non-monotonic behavior of chemical potential as a function of Gauss-Bonnet coupling at different temperatures. At an elevated temperature related to the left panel of Fig. 2, the chemical potential decreases as the coupling constant increases. Conversely, at a lower temperature corresponding to the middle panel of Fig. 2, the chemical potential increases monotonically with increasing coupling. The above results are explicitly shown in the top-left and top-right panels of Fig. 5.
@@
Figure 6: The free energy and its first-order derivative w.r.t. the chemical potential as functions of temperature for different chemical potentials. The blue curves represent the holographic superconductors while the red curves describe the spontaneous scalarization, and the vertical gray dashed line indicates the transition point.
@@ III.3 Instability analysis using effective potential
In this subsection, we complement our analysis by studying the tachyonic instability of the effective potentials. On the one hand, for a holographic superconductor, it is understood that the coupling between the scalar and Maxwell field in asymptotic AdS spacetime leads to tachyonic instability [17]. On the other hand, the coupling between the scalar field and the Gauss-Bonnet curvature also gives rise to tachyonic instability [20, 49, 50], and it has been argued that the origin of the instability is Gregory-Laflamme type [48]. This section delves into the effective potential Eq. (11) of the scalar perturbations regarding the underlying instabilities. In literature, the tachyonic instability is primarily attributed to a negative effective mass extracted from examining the master equation of scalar perturbations. For a holographic superconductor, a negative effective scalar mass arises from the non-vanishing the Maxwell field (electrostatic potential) in AdS spacetime. For spontaneous scalarization, this is due to the contribution coming from the non-vanishing coupling between the scalar and higher curvature term. In what follows, the numerical results of the effective potential Eq. (11) are obtained by solving the system of equations, Eqs. (8) and (9), for given $\lambda$, $q$, and $\phi^{\prime}(r_{h})=0.5$ instead of $\psi(r_{h})$.
@@
In Fig. 8, we present the effective potentials and the corresponding profiles of the scalar field, evaluated for various metric parameters. As pointed out in [48], the effective potential of tachyonic instability in AdS spacetime is featured by a positive barrier near the event horizon that smoothly converges to a given value at infinity. In the top-left panel of Fig. 8, the resulting effective potentials of a pure holographic superconductor agree with such a behavior. In the present model, this can be achieved by assuming $\lambda=0$. We note that the effective potential asymptotically approaches $V_{\mathrm{eff}}\to-q^{2}\phi(r)^{2}$ at spatial infinity for the given metric parameters. The different curves in the plot are obtained by taking different values for the charge $q$. A positive potential barrier is formed near the event horizon for all the cases. An increase of the charge $q$ causes the potential barrier to dissipate less rapidly as the radial coordinate increases. The negative values of the effective potential are attributed to the condensation of the Maxwell field, giving rise to tachyonic instability in scalar perturbations. Also, a more significant coupling between the scalar and Maxwell fields leads to a higher transition temperature for holographic superconductivity [17].
@@
Figure 8: The top row shows the calculated effective potentials for different model parameters. Top-left: The effective potential for the holographic superconductors with $\lambda=0$, evaluated for different charges $q$. Top-right: The effective potential for the spontaneously scalarized black hole with $q=0$, evaluated for different values of $\lambda$. The bottom row shows the corresponding profiles of the scalar field $\psi(r)$.
@@
Conversely, a pure spontaneously scalarized black hole metric can be obtained in the present model by taking $q=0$. The corresponding effective potentials are evaluated and shown in the top-right panel of Fig. 8, where one varies the Gauss-Bonnet coupling $\lambda$. One observes that the obtained effective potential’s main feature differs from that of a holographic superconductor. In particular, a potential well is formed near the event horizon, and the depth of this well gradually increases as the Gauss-Bonnet coupling $\lambda$ grows. Therefore, the potential well close to the horizon caused by the Gauss-Bonnet coupling is primarily understood to cause the tachyonic instability. Such a characteristic of the effective potential has been extensively discussed in the literature of spontaneous scalarization [20, 47], which was attributed to the Gregory-Laflamme instability in Ref. [48]. The above discussions further confirm the distinct nature of the holographic superconductivity and spontaneous scalarization phases, which are substantially triggered by entirely distinct mechanisms. Besides the thermodynamic quantities, such as the free energy, the difference is also manifested by the effective potential.